# GnnMerge: Merging of Gnn Models Without Accessing Training Data

### v1

@@ 1 Introduction
Given two neural models, can we merge them into a single model integrating the capabilities of both, without accessing the original training data? This is the core question driving the emergent field of model merging (Stoica et al., 2024; Ainsworth et al., 2023; Yang et al., 2024b; Ilharco et al., 2023; Huang et al., 2024; Lu et al., 2024). Model merging addresses key challenges in dynamic machine learning environments where retraining-from-scratch is impractical or impossible. For instance, the introduction of training data annotated with new class labels—such as novel research areas in citation networks or new product types in e-commerce— necessitates full retraining after incorporating the new training data. A more efficient alternative would be to train a new model exclusively on the new data and then merge it with the existing model, and thereby eliminating the need for full retraining. Similarly, one may wish to merge two models trained on the same dataset but for different tasks into a single multi-task model. In privacy-sensitive settings, organizations may wish to combine independently trained models without sharing raw data, avoiding privacy breaches or exposing proprietary information.
@@ 1.1 Existing works and Limitations
Table 1: Characterization of existing algorithms for model merging: denotes the presence of an undesirable property, whereas indicates its absence. While a numerical learning-based optimization can be an effective model merging procedure, it also results in higher computational costs. In this context, a specifically highlights this increased computational burden.
@@
  • Model inflation: The number of parameters in a model directly impacts its computational efficiency and GPU memory requirements. Ideally, the merged model should maintain the same size as the individual models being merged to preserve these efficiencies. However, several existing algorithms fail to meet this desideratum, leading to inflated model size with increased resource demands. Inflation may happen due to various design choices, such as the injection of adapter layers between model layers (Yang et al., 2024a) or partial merging of layers to avoid degradation of performance (Stoica et al., 2024).
  • Numerical learning-based merging: Merging algorithms can broadly be divided into two categories. The first category employs analytical operations on the input model parameters to produce the merged model. The second category adopts a numerical learning-based approach, optimizing the merged model’s parameters by minimizing a loss function. While this method achieves better accuracy, it compromises on computational efficiency. Given that Gnns predominantly use a message-passing paradigm (Kipf & Welling, 2017; Xu et al., 2019; Veličković et al., 2018; Hamilton et al., 2017b), can we achieve both efficiency and efficacy?
@@ 2 Problem Formulation
For node classification or regression, the components in $\mathbb{T}$ are nodes, while for link prediction, they correspond to edges. Similarly, the ground-truth labels in $\mathcal{Y}$ indicate class labels for node classification and the presence or absence of edges for link prediction. Commonly used loss functions include cross-entropy, negative log-likelihood, and RMSE.
The problem of model merging is now defined as follows.
@@ 3.2 Merging through Node Embedding Alignment
$\overbrace{\sum_{i=1}^{n}}^{\begin{subarray}{c}\text{For each}\\ \text{base model,}\end{subarray}}\overbrace{\sum_{\ell=1}^{L}}^{\begin{subarray}{c}\text{{Gnn}}\\ \text{layer}\end{subarray}}\overbrace{\sum_{\forall v\in\mathcal{V}}}^{\text{node}}$ $\|\Theta_{M}^{\ell}\left(\mathbf{h}_{v,M}^{\ell-1},\mathcal{S}_{v,M}^{\ell-1}\right)$ (5)
$-\Theta_{i}^{\ell}\left(\mathbf{h}_{v,i}^{\ell-1},\mathcal{S}_{v,i}^{\ell-1}\right)\|_{2}$
Here, $\mathbf{h}_{v,i}^{\ell-1}$ represents the embedding of node $v$ and $\mathcal{S}_{v,i}^{\ell-1}$ denotes the embeddings of its neighbors in model $\Theta_{i}$ from layer $\ell-1$ (analogously defined for $\Theta_{M}$). The parameters $\Theta_{i}^{\ell}$ are responsible for the transformations at layer $\ell$. Our objective is to determine the parameters of $\Theta_{M}^{\ell}$ for each layer $\ell$, ensuring that the merged model $\Theta_{M}$ generates embeddings that closely align with those produced by the base models.
Although the minimization task described in Eq. 5 is both task-agnostic and independent of training labels, it is computationally intensive. The process is equivalent to training a student Gnn ($\Theta_{M}$) to mimic a set of teacher Gnns (the base models). This approach contradicts the computational efficiency requirements outlined in § 2. However, we will demonstrate how a minor relaxation of the formulation can dramatically reduce the computational burden, aligning with our efficiency goals.
@@ 3.3 Independent Node Embedding Alignment
Fig. 2 provides a visual comparison between the original and the relaxed objectives. The key insight behind this relaxation is that the learning objective for $\Theta_{M}^{\ell}$ becomes independent of the preceding layers of the merged model since its input is no longer derived from the outputs of its own previous layers. Intuitively, this adjustment directs each layer of the merged model toward a parameter space where the linear transformations applied to the node embeddings of the base models (rather than its own embeddings) closely approximate the transformations induced by the base model’s parameters.
Figure 2: Inputs according to Joint Embedding Alignment (Eq. 5) [left] and the Independent Embedding Alignment (Eq. 6) [right].
This relaxation is expected to have a mild effect since, at layer 00, all models share the same input, i.e., $\mathbf{h}^{0}_{v}=\mathbf{x}_{v}$. Consequently, for a 1-layer Gnn, the relaxed objective in Eq. 6 is equivalent to the original objective in Eq. 5. For deeper Gnns, if $\Theta_{M}^{1}$ effectively approximates the base models, then $\mathbf{h}_{v,M}^{\ell}\approx\mathbf{h}_{v,i}^{\ell}$ and $\mathbf{S}_{v,M}^{\ell}\approx\mathbf{S}_{v,i}^{\ell}$, resulting in transitive consistency across subsequent layers. We further note that Gnns are typically not deep (often $\leq 3$ layers) due to the well-established problems of oversquashing and oversmoothing (Rusch et al., 2023; Giovanni et al., 2024).
Next, we demonstrate how this relaxation enables analytical solutions for popular Gnn architectures, resulting in dramatic efficiency improvements.
@@ 3.4 Analytical Solution
were $\|A\|^{2}_{F}$ represents the frobenius norm of matrix $A$. Since Eq. 9 is convex, the minima is achieved when the gradient with respect to $\mathbf{W}_{k,M}^{\ell}$ is 00. Setting it to 0, we get:
$\sum_{i=1}^{n}\left[2(\mathbf{W}_{k,M}^{\ell})^{\mathtt{T}}(\mathbf{Z}_{k,i}^{\ell-1})^{\mathtt{T}}\mathbf{Z}_{k,i}^{\ell-1}-2(\mathbf{G}_{K,i}^{\ell})^{\mathtt{T}}\mathbf{G}_{K,i}^{\ell}\right]=0$
or, $(\mathbf{W}_{k,M}^{\ell})^{\mathtt{T}}\sum_{i=1}^{n}(\mathbf{Z}_{k,i}^{\ell-1})^{\mathtt{T}}\mathbf{Z}_{k,i}^{\ell-1}=\sum_{i=1}^{n}(\mathbf{G}_{K,i}^{\ell})^{\mathtt{T}}\mathbf{G}_{K,i}^{\ell}$
or, $(\mathbf{W}_{k,M}^{\ell})^{\mathtt{T}}=\sum_{i=1}^{n}(\mathbf{G}_{K,i}^{\ell})^{\mathtt{T}}\mathbf{G}_{K,i}^{\ell}(\sum_{i=1}^{n}(\mathbf{Z}_{k,i}^{\ell-1})^{\mathtt{T}}\mathbf{Z}_{k,i}^{\ell-1})^{-1}$ (10)
@@ 4 Experiments
The implementation of our algorithm is available at https://github.com/idea-iitd/GNNMerge/tree/main.
@@ 4.1 Experimental Setup
    Node Classification on In-domain Datasets: We create two disjoint label splits from the same dataset and train a model on each split independently. The merging process is then performed on these two models, simulating a scenario where new labels are introduced after the initial training phase.
@@
    Node Classification on Different Datasets: Given $N$ models, each trained for node classification on a distinct dataset, we merge these models into a single unified model. The performance of the merged model is subsequently evaluated on the test sets of the respective datasets. To ensure a common architecture for Gnn models across different datasets, we only the utilize Text-Attributed Graphs (TAGs) from Table 3 (first five rows). Raw text attributes associated with nodes in these datasets are processed using Sentence-BERT (Reimers, 2019) to generate uniform 384-dimensional feature representations for all datasets.
@@
Datasets. Table 3 lists the $8$ graph datasets used for our experiments. They span graphs of various sizes, densities, and class label sizes.
Dataset #Nodes #Edges #Classes #Features
Cora (Yang et al., 2016) 2,708 5,429 7 1,433
Citeseer (Yang et al., 2016) 3,312 4,732 6 3,703
Pubmed (Yang et al., 2016) 19,717 44,338 3 500
Arxiv (Hu et al., 2020) 169,443 2,315,598 40 128
WikiCS (Mernyei & Cangea, 2020) 11,701 431,726 10 300
AmzPhoto (Shchur et al., 2018) 7,650 238,162 8 745
AmzComp (Shchur et al., 2018) 13,752 491,722 10 767
Reddit (Hamilton et al., 2017a) 232,965 114,615,892 41 602
Table 3: Datasets used for benchmarking GnnMerge.
Baselines. All the existing algorithms listed in Table 1 with a in the “Inhibiting properties” column are inapplicable in our setting. These include methods requiring labeled data (e.g., Fisher Merging, UQ-Merge) or those dependent on a common pre-trained backbone fine-tuned for all tasks (e.g., Task Arithmetic, AdaMerging, etc.). After excluding such algorithms, we focus on the remaining applicable methods: Weight Averaging (WAvg.), Git-Rebasin, Permute, and ZipIt!. Additionally, we compare against Surgery, a post-hoc refinement method applied to a merged model. Surgery supports a variant where WAvg. is used to create the merged model, making it compatible with our setting. We use GnnMerge to denote the layer-independent learning-based methodology proposed in this work (§ 3.3) and GnnMerge++ to denote the analytical version (§ 3.4).
@@ 4.2 Results
Tasks Raw Random WAvg. Permute ZipIt! Surgery GnnMerge GnnMerge++
Arxiv-NC 73.10 9.31 57.99 61.03 60.72 66.54 73.03 73.01
Pubmed-LP 97.05 90.22 91.32 91.78 91.50 93.67 96.16 96.23
WikiCS-NC 79.32 13.99 75.18 74.79 75.88 77.02 78.79 78.88
Cora-LP 94.34 83.62 77.28 76.90 80.56 88.39 94.12 94.45
WikiCS-NC 79.32 22.87 69.46 71.30 73.35 76.93 78.88 78.91
Pubmed-LP 97.05 91.50 88.70 90.07 89.45 92.71 96.26 96.37
[…]
Different Datasets. Table 8 presents the performance of merging Gcns trained for node classification across two distinct datasets. Both GnnMerge and GnnMerge++ achieve average accuracies that are comparable to the base models. Additionally, GnnMerge outperforms the closest baseline by 7.02%. While Surgery is competitive in some cases, it must be noted that it comes at the cost of model inflation by introducing task-specific parameters.
To further stress-test the methods, we extend the analysis by merging more than two models trained on multiple datasets. The full results of this evaluation are provided in Tables 910, and 11 in the appendix. In Fig. 3, we present the average accuracy of the merged models across all dataset combinations of a particular size (i.e., the row corresponding to “Average” in Tables 8,  910, and 11). In this analysis, we have excluded Git-Rebasin since the source code does not support merging of more than two models, and its performance is not competitive even for two datasets (Table 8). Additionally, GnnMerge++ is omitted from Fig. 3 since its performance closely mirrors that of GnnMerge (see Tables 8910, and 11 in the Appendix).
As depicted in Fig. 3, GnnMerge demonstrates significantly superior robustness when merging multiple models, exhibiting only a negligible decline in performance even when merging up to five models. Notably, when merging five models, GnnMerge achieves an impressive 23.53% improvement over the best baseline.
Different Tasks. Table 5 presents results for a more challenging scenario: merging Gcns trained on two distinct tasks—node classification and link prediction—across two different datasets. In most cases, baseline methods perform no better than a randomly initialized Gcn on the link prediction task. Preserving node classification accuracy leads to a severe degradation in link prediction AUC for the baselines. In contrast, GnnMerge and GnnMerge++ achieve accuracies comparable to the individual models on their respective tasks.
@@
4.3 Ablation Study
We aim to address two key questions in the next experiment. First, how does the number of Gnn layers impact model merging performance? Second, as discussed in § 3.3, learning parameters through joint node embedding alignment introduces computational overhead and slower convergence (Eq. 5). To mitigate this, we propose layer-wise independent node embedding alignment, which serves as a relaxation of the original objective (Eq. 6). What effect does this relaxation have on performance?
(a) Model 1:Citeseer, Model 2:Pubmed
(b) Overlap with GnnMerge++
(c) Overlap with WAvg.
Arch. Datasets Raw WAvg. Permute ZipIt! Surgery GnnMerge GnnMerge++
GraphSage Arxiv 74.65 55.13 58.45 59.13 68.39 72.85 72.78
WikiCS 78.82 72.84 56.17 56.68 69.22 78.65 78.63
Arxiv 74.65 68.77 60.00 63.60 67.85 74.24 74.29
Pubmed 77.96 72.32 62.21 65.94 75.12 78.01 77.97
NodeFormer Cora 81.09 50.72 - - 77.61 76.59 75.27
Citeseer 81.35 71.71 - - 79.15 77.83 76.08
Pubmed 80.08 57.60 - - 75.74 79.78 79.10
WikiCS 74.17 62.80 - - 67.48 72.42 69.21
[…]
Fig. 4 presents the performance of the two optimization strategies as we vary the number of Gcn layers in the merging models on the arXiv dataset. A clear trend emerges: as the number of Gnn layers increases, the performance of the joint node alignment strategy deteriorates (indicated by a growing gap with the base model accuracies). In contrast, the relaxed optimization strategy, which aligns each layer independently, remains stable and does not suffer from this degradation. This behavior is attributed to the vanishing gradient problem becoming more pronounced in joint node alignment as the number of layers increases. Treating layers independently circumvents this issue, as the optimization problem remains decoupled from the depth of the network.
4.4 Generalization to Gnn Architectures
We further benchmark GnnMerge on GraphSage and NodeFormer in Tables 6 and 12 (in Appendix). The results follow the same trend observed with Gcn on both architectures, and thereby establishing the robustness of GnnMerge to accommodate diverse Gnn architectures.
4.5 Efficiency
Dataset Scratch Train Time GnnMerge GnnMerge++
Arxiv 24.19s 3.75s 1.67s
Reddit 697.88s 102.99s 5.12s
Table 7: Running times for the In-domain dataset task.
@@ 5 Conclusions
Statement of Impact
This work aims to advance the field of Machine Learning by addressing the unexplored challenge of model merging in Graph Neural Networks (Gnns). By proposing a novel, task-agnostic node embedding alignment strategy and introducing computationally efficient analytical solutions for widely-used GNN architectures, our approach has the potential to benefit real-world applications that rely on GNNs, such as social network analysis, recommendation systems, and biological graph modeling.
The societal implications of this work include enabling more efficient use of computational resources, as our method reduces the need for retraining models from scratch, which can save time, energy, and cost. Moreover, this efficiency may facilitate the use of GNNs in resource-constrained environments. We do not foresee any ethical concerns arising out of our work.

### v2

@@ 1 Introduction
Given two neural models, can we merge them into a single model integrating the capabilities of both, without accessing the original training data? This is the core question driving the emergent field of model merging Stoica et al. (2024); Ainsworth et al. (2023); Yang et al. (2024b); Ilharco et al. (2023); Huang et al. (2024); Lu et al. (2024). Model merging addresses key challenges in dynamic machine learning environments where retraining-from-scratch is impractical or impossible. For instance, the introduction of training data annotated with new class labels—such as novel research areas in citation networks or new product types in e-commerce— necessitates full retraining after incorporating the new training data. A more efficient alternative would be to train a new model exclusively on the new data and then merge it with the existing model, and thereby eliminating the need for full retraining. Similarly, one may wish to merge two models trained on the same dataset but for different tasks into a single multi-task model. In privacy-sensitive settings, organizations may wish to combine independently trained models without sharing raw data, avoiding privacy breaches or exposing proprietary information.
@@ 1.1 Existing works and Limitations
Table 1: Characterization of existing algorithms for model merging: denotes the presence of an undesirable property, whereas ✗ indicates its absence. While a numerical learning-based optimization can be an effective model merging procedure, it also results in higher computational costs. In this context, a specifically highlights this increased computational burden.
@@
  • Model inflation: The number of parameters in a model directly impacts its computational efficiency and GPU memory requirements. Ideally, the merged model should maintain the same size as the individual models being merged to preserve these efficiencies. However, several existing algorithms fail to meet this desideratum, leading to inflated model size with increased resource demands. Inflation may happen due to various design choices, such as the injection of adapter layers between model layers Yang et al. (2024a) or partial merging of layers to avoid degradation of performance Stoica et al. (2024).
  • Numerical learning-based merging: Merging algorithms can broadly be divided into two categories. The first category employs analytical operations on the input model parameters to produce the merged model. The second category adopts a numerical learning-based approach, optimizing the merged model’s parameters by minimizing a loss function. While this method achieves better accuracy, it compromises on computational efficiency.
@@ 2 Problem Formulation
For node classification or regression, the components in $\mathbb{T}$ are nodes, while for link prediction, they correspond to edges. Similarly, the ground-truth labels in $\mathcal{Y}$ indicate class labels for node classification and the presence or absence of edges for link prediction. Commonly used loss functions include cross-entropy, negative log-likelihood, and RMSE. The problem of model merging is now defined as follows.
@@ 3.2 Merging through Node Embedding Alignment
$\overbrace{\sum_{i=1}^{n}}^{\begin{subarray}{c}\text{For each}\\ \text{base model,}\end{subarray}}\overbrace{\sum_{\ell=1}^{L}}^{\begin{subarray}{c}\text{{Gnn}}\\ \text{layer}\end{subarray}}\overbrace{\sum_{\forall v\in\mathcal{V}}}^{\text{node}}$ $\|\Theta_{M}^{\ell}\left(\mathbf{h}_{v,M}^{\ell-1},\mathcal{S}_{v,M}^{\ell-1}\right)-\Theta_{i}^{\ell}\left(\mathbf{h}_{v,i}^{\ell-1},\mathcal{S}_{v,i}^{\ell-1}\right)\|_{2}$ (5)
Here, $\mathbf{h}_{v,i}^{\ell-1}$ represents the embedding of node $v$ and $\mathcal{S}_{v,i}^{\ell-1}$ denotes the embeddings of its neighbors in model $\Theta_{i}$ from layer $\ell-1$ (analogously defined for $\Theta_{M}$). The parameters $\Theta_{i}^{\ell}$ are responsible for the transformations at layer $\ell$. Our objective is to determine the parameters of $\Theta_{M}^{\ell}$ for each layer $\ell$, ensuring that the merged model $\Theta_{M}$ generates embeddings that closely align with those produced by the base models. Although the minimization task described in Eq. 5 is both task-agnostic and independent of training labels, it is computationally intensive. The process is equivalent to training a student Gnn ($\Theta_{M}$) to mimic a set of teacher Gnns (the base models). This approach contradicts the computational efficiency requirements outlined in § 2. However, we will demonstrate how a minor relaxation of the formulation can dramatically reduce the computational burden, aligning with our efficiency goals.
@@ 3.3 Independent Node Embedding Alignment
The key insight behind this relaxation is that the learning objective for $\Theta_{M}^{\ell}$ becomes independent of the preceding layers of the merged model since its input is no longer derived from the outputs of its own previous layers. Intuitively, this adjustment directs each layer of the merged model toward a parameter space where the linear transformations applied to the node embeddings of the base models (rather than its own embeddings) closely approximate the transformations induced by the base model’s parameters. This relaxation is expected to have a mild effect since, at layer 00, all models share the same input, i.e., $\mathbf{h}^{0}_{v}=\mathbf{x}_{v}$. Consequently, for a 1-layer Gnn, the relaxed objective in Eq. 6 is equivalent to the original objective in Eq. 5. For deeper Gnns, if $\Theta_{M}^{1}$ effectively approximates the base models, then $\mathbf{h}_{v,M}^{\ell}\approx\mathbf{h}_{v,i}^{\ell}$ and $\mathbf{S}_{v,M}^{\ell}\approx\mathbf{S}_{v,i}^{\ell}$, resulting in transitive consistency across subsequent layers. We further note that Gnns are typically not deep (often $\leq 3$ layers) due to the well-established problems of oversquashing and oversmoothing (Rusch et al., 2023; Giovanni et al., 2024). Next, we demonstrate how this relaxation enables analytical solutions for popular Gnn architectures, resulting in dramatic efficiency improvements.
@@ 3.4 Analytical Solution
were $\|A\|^{2}_{F}$ represents the frobenius norm of matrix $A$. Since Eq. 9 is convex, the minima is achieved when the gradient with respect to $\mathbf{W}_{k,M}^{\ell}$ is 00. Setting it to 0 and solving for $\mathbf{W}_{k,M}^{\ell}$, we get:
$(\mathbf{W}_{k,M}^{\ell})^{\mathtt{T}}=\sum_{i=1}^{n}(\mathbf{G}_{K,i}^{\ell})^{\mathtt{T}}\mathbf{G}_{K,i}^{\ell}(\sum_{i=1}^{n}(\mathbf{Z}_{k,i}^{\ell-1})^{\mathtt{T}}\mathbf{Z}_{k,i}^{\ell-1})^{-1}$ (10)
@@ 4 Experiments
The implementation of our algorithm is available at https://anonymous.4open.science/r/Model-Merging-GNNs-4C55.
@@ 4.1 Experimental Setup
    Node Classification on In-domain Datasets: We create two disjoint label splits from the same dataset and train a model on each split independently. The merging process is then performed on these two models, simulating a scenario where new labels are introduced after the initial training.
@@
    Node Classification on Different Datasets: Given $N$ models, each trained for node classification on a distinct dataset, we merge these models into a single unified model. The performance of the merged model is subsequently evaluated on the test sets of the respective datasets. To ensure a common architecture for Gnn models across different datasets, we only the utilize Text-Attributed Graphs from Table 4 (first five rows). Raw text attributes associated with nodes in these datasets are processed using Sentence-BERT (Reimers, 2019) to generate uniform feature representations.
@@
Datasets. Table 4 lists the $8$ graph datasets used for our experiments.
Baselines. All the existing algorithms listed in Table 1 with a in the “Inhibiting properties” column are inapplicable in our setting. These include methods requiring labeled data (e.g., Fisher Merging, UQ-Merge) or those dependent on a common pre-trained backbone fine-tuned for all tasks (e.g., Task Arithmetic, AdaMerging, etc.). After excluding such algorithms, we focus on the remaining applicable methods: Weight Averaging (WAvg.), Git-Rebasin, Permute, and ZipIt!. Additionally, we compare against Surgery, a post-hoc refinement method applied to a merged model. Surgery supports a variant where WAvg. is used to create the merged model, making it compatible with our setting. We use GnnMerge to denote the layer-independent learning-based methodology proposed in this work and GnnMerge++ to denote the analytical version.
@@ 4.2 Results
Different Datasets. Table 9 presents the performance of merging Gcns trained for node classification across two distinct datasets. Both GnnMerge and GnnMerge++ achieve average accuracies that are comparable to the base models. Additionally, GnnMerge outperforms the closest baseline by 7.02%. While Surgery is competitive in some cases, it must be noted that it comes at the cost of model inflation by introducing task-specific parameters. To further stress-test the methods, we extend the analysis by merging more than two models trained on multiple datasets. The full results of this evaluation are provided in Tables 1011, and 12 in the appendix. In Fig. 4, we present the average accuracy of the merged models across all dataset combinations of a particular size (i.e., the row corresponding to “Average” in Tables 9,  1011, and 12). In this analysis, we have excluded Git-Rebasin since the source code does not support merging of more than two models, and its performance is not competitive even for two datasets (Table 9). Additionally, GnnMerge++ is omitted from Fig. 4 since its performance closely mirrors that of GnnMerge (see Tables 91011, and 12 in the Appendix). As depicted in Fig. 4, GnnMerge demonstrates significantly superior robustness when merging multiple models, exhibiting only a negligible decline in performance even when merging up to five models. Notably, when merging five models, GnnMerge achieves an impressive 23.53% improvement over the best baseline.
Results on generalization to different architectures and merging of different tasks are present in Appendix Section C.
4.3 Ablation Study
@@
We aim to address two key questions in the next experiment. First, how does the number of Gnn layers impact model merging performance? Second, as discussed in § 3.3, learning parameters through joint node embedding alignment introduces computational overhead and slower convergence (Eq. 5). To mitigate this, we propose layer-wise independent node embedding alignment, which serves as a relaxation of the original objective (Eq. 6). What effect does this relaxation have on performance? Fig. 2 presents the performance of the two optimization strategies as we vary the number of Gcn layers in the merging models on the arXiv dataset. A clear trend emerges: as the number of Gnn layers increases, the performance of the joint node alignment strategy deteriorates. In contrast, the relaxed optimization strategy, which aligns each layer independently, remains stable and does not suffer from this degradation. This behavior is attributed to the vanishing gradient problem becoming more pronounced in joint node alignment as the number of layers increases. Treating layers independently circumvents this issue, as the optimization problem remains decoupled from the depth of the network.
Figure 3: Visual Illustration of embedding alignment using GnnMerge and WAvg.